a rectangular tank of length 16cm,width 9 cm and height 13cm contain water upto a height of 8 cm.calculate the total surface area of the tank that is in contact with the water?

Answer by Cromlix(4381)

The Surface Area in contact with
the water would be:-
2 * 8cm * 16cm = 256cm^2 (long sides)
1 * 16cm * 9cm = 144cm^2 (base)
2 * 8cm * 9cm = 144cm^2 (short sides)
TOTAL = 544cm^2.
Hope this helps.
:-)